OpenStudy (anonymous):

2. Solve the compound inequality 3x + 4 > 5 and x – 2 < 7. Clearly show and explain each step ??

OpenStudy (anonymous):

Solve each inequality separately: 3x+4>5 and x-2<7 3x>1 subtracting 4 from both sides. x>1/3 divide both sides by 3. Can you solve for x in the second inequality and tell me what you got?

OpenStudy (anonymous):

i have no clue what you just did.. so i cant..

OpenStudy (anonymous):

Solve each inequality for x, that is what I did.

OpenStudy (anonymous):

treat inequality exactly as an equation, except when you multiply or divide by a negative number, which we are not doing here. So look at the statements again. and solve both separately to find the value of x

OpenStudy (anonymous):

You can do the same operations as you would when doing a linear equation. 3x+4>5 3x>1 subtracting 4 from both sides. x>1/3 divide both sides by 3. Do you not understand this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then solve for x: x-2<7. What do you get?

OpenStudy (anonymous):

x-2<7 x<7 add 2 to both sides. ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes. So you should have: x<9

OpenStudy (anonymous):

oh yeah lol.. so whats the final answer for the problem?

OpenStudy (anonymous):

So after solving you have: x>1/3 and x<9, we need to figure what range of values for x in both cases would make both equations true. Is x=2 a solution. That is 2>1/3 and 2<9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the values for x that would be a solution has to fall within 1/3 and 9: Answer: {x| x>1/3 and x<9} or {x|1/3<x<9}

OpenStudy (anonymous):

Anymore questions?

OpenStudy (anonymous):

3. Solve = 5. ? what does this mean?

OpenStudy (anonymous):

3. Solve l4x-3l = 5

OpenStudy (anonymous):

@philo1234 ?

OpenStudy (anonymous):

Solving absolute equation: 1. Ensure that the absolute expression is by itself on one side |4x-3|=5 2. Check: The absolute value of an expression has to be equal to or greater than 0. If negative then there is no solution otherwise there is a solution. 2. 2 cases: Case1: Case 2: 4x-3 = 5 or 4x-3= -5 Solve for x in both cases. Can you do them and tell me what you got?

OpenStudy (anonymous):

@ethompsonn ?

OpenStudy (anonymous):

sorry i got it nvm thx tho

OpenStudy (anonymous):

k